OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, -3) and has a slope of 1/2 ? 2x + y = 4 2x - y = 4 x + 2y = 4 x - 2y = 4

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@ybarrap can you please help me

OpenStudy (ybarrap):

you have the slope and so we just need to find the y-intercept: y=mx+b, where m is the slope (1/2) an b is the y-intercept Take the point (-2,-3) and just plug into y=mx+b to find b: -3=(1/2) * (-2) + b b=-3 +1=-2 So, y = (1/2) x - 2 Multiply both sides by 2: 2y - x = -4 Multiply both sides by -1: x - 2y = 4 That's it.

OpenStudy (anonymous):

thank you so much!! (: